T/F/U. Imagine that two goods make you happy: Good A and good B. The marginal utility of good A is 10 at a price of $2 per unit and the marginal utility of good B is 5 at a price of $1per unit. To satisfy the optimal consumption rule, you would want twice as many of good B than good A. Show your work and tell me how you arrived at your answer using the definition of optimal consumption.
Answer #1

False

MRS= MUa/MUb = 10/5 = 2

Pa/Pb= 2/1= 2

At optimum MRS= Pa/Pb

This condition has been satisfied. So the consumer will not change his combination of goods.